### v2

# SeCon-RAG: A Two-Stage Semantic Filtering and Conflict-Free Framework for Trustworthy RAG

Abstract
Retrieval-augmented generation (RAG) systems enhance large language models (LLMs) with external knowledge but are vulnerable to corpus poisoning and contamination attacks, which can compromise output integrity. Existing defenses often apply aggressive filtering, leading to unnecessary loss of valuable information and reduced reliability in generation. To address this problem, we propose a two-stage semantic filtering and conflict-free framework for trustworthy RAG. In the first stage, we perform a joint filter with semantic and cluster-based filtering which is guided by the Entity-intent-relation extractor (EIRE). EIRE extracts entities, latent objectives, and entity relations from both the user query and filtered documents, scores their semantic relevance, and selectively adds valuable documents into the clean retrieval database. In the second stage, we proposed an EIRE-guided conflict-aware filtering module, which analyzes semantic consistency between the query, candidate answers, and retrieved knowledge before final answer generation, filtering out internal and external contradictions that could mislead the model. Through this two-stage process, SeCon-RAG effectively preserves useful knowledge while mitigating conflict contamination, achieving significant improvements in both generation robustness and output trustworthiness. Extensive experiments across various LLMs and datasets demonstrate that the proposed SeCon-RAG markedly outperforms state-of-the-art defense methods.
1 Introduction
Large Language Models (LLMs) (bai2023qwen, ; duan2025oyster, ; chatgpt, ) have demonstrated remarkable capabilities across a wide range of natural language tasks (zhao2023survey, ; fang2025hierarchical, ; fang2025turing, ). However, they still suffer from critical security vulnerabilities, including adversarial attacks (yu2025infecting, ; mckenzie2025stack, ), jailbreak attacks (jia2024improved, ; huang2024semantic, ; yang2025cannot, ), and other alignment challenges. Moreover, their knowledge is fundamentally limited by their training data, which can lead to outdated or hallucinated information. Retrieval-Augmented Generation (RAG) addresses this issue by dynamically incorporating external documents during generation, improving factual accuracy and timeliness (lewis2020retrieval, ; arslan2024survey, ). However, due to the reliance on external corpora, RAG systems are susceptible to corpus poisoning and retrieval contamination attacks, which involve injecting adversarial content into the retrieval database to manipulate the model’s output (nazary2025poison, ; zhang2025practical, ; chang2025one, ).
Recent defense strategies have attempted to address this by employing adversarial training, retrieval filtering and reasoning-based conflict resolution (xiang2024certifiably, ; wang2024astute, ; zhou2025trustrag, ). These methods primarily use Coarse-grained filtering or voting to remove malicious documents, and the inference phase does not consider what information the RAG should select when confronted with conflicting content, which can result in two limitations. (1) Coarse-Grained filtering will removes both harmful and useful content. (2) Failure to resolve conflicts between retrieved and the LLM’s internal knowledge, which leads to untrustworthy results.
To address these issues, our framework first integrates semantic information into the RAG filtering method. We extract intrinsic semantic signals from each document to allow for fine-grained filtering while also facilitating the resolution of conflicting evidence during inference. Building on this insight, we propose SeCon-RAG, a two-stage framework that combines semantic and cluster-based filtering with conflict-filtering retrieval-augmented generation.
We first design a semantic extraction module called EIRE (Entity-Intent-Relation Extractor). It makes future modules easier to use by extracting entities, hidden intentions, and relationships between entities from document information. In the first stage, we propose a Semantic and Clustering-Based Filtering module (SCF) based on EIRE. On the one hand, it filters the intensive incorrect documents based on their cluster in the embedding space. On the other hand, using EIRE, the semantic structure graph of candidate documents and verified correct documents can help to exclude more hidden poisoned documents. The implementation of this dual filtering mechanism can ensure that the majority of malicious and poisonous documents are filtered out while also preventing potentially valuable documents from being wasted.
In the second stage, we propose an EIRE-guided conflict-aware filtering (CAF) module that checks the semantic consistency of the query, the candidate context, and the model’s internal knowledge. CAF uses EIRE to extract semantic information from the final input information, judge different information based on semantic knowledge, and remove misleading information caused by internal and external knowledge conflicts or omissions before generating the final response. This ensures that the final generations are not only factually accurate, but also semantically consistent across internal and external knowledge sources.
In comparison to previous work, our work makes significant advances. Our approach is the first to incorporate semantic information into the retrieval and inference phases of RAG defenses. We propose a two-stage defense framework that employs semantic reasoning to ensure robust during retrieval (SCF) and generation (CAF). The proposed framework implements structured semantic filtering by extracting entity-intent relationships and using them to filter poisoning documents which may evade clustering-based defenses.
We evaluate SeCon-RAG on three challenging QA benchmarks Natural Questions, HotpotQA, and MS-MARCO across five different LLMs including LLaMA-3.1-8B (dubey2024llama, ), Mistral-12B (mistral_nemo_instruct_2407, ), GPT-4o (achiam2023gpt, ), DeepSeek-R1 (guo2025deepseek, ), and Qwen-7B (hui2024qwen2, ). Our method consistently improves robustness, consistency, and resistance to corpus poisoning across all settings. Our main contributions are summarized as follows:
(a) We are the first to incorporate structured semantic information into RAG defense filtering by the proposed EIRE module, allowing for fine-grained understanding of entity, intent, and relation structures to improve the precision of poisoned content detection. (b) Building on EIRE, we propose SeConRAG, a two-stage defense framework that combines semantic and cluster-based filtering with conflict-aware filtering to improve retrieval robustness and answer consistency. (c) Extensive experiments on a variety of datasets and LLMs show that SeConRAG consistently achieves high factual accuracy, low attack success rates, and high generalizability, demonstrating its practical effectiveness and plug-and-play capabilities.
2 Related Works
2.1 Retrieval-Augmented Generation
Retrieval-Augmented Generation improves large language models by supplementing them with external knowledge extracted from large corpora, thereby addressing limitations in factual recall and knowledge coverage (lewis2020retrieval, ; wu2024retrieval, ). While RAG’s generation quality has improved, it continues to suffer from retrieval errors, hallucinations, and poor content integration. To address these issues, previous research has focused on query rewriting, index optimization, and memory-based retrieval (zheng2023take, ; ma2023query, ). Recent LLM-augmented methods include Insight-RAG (pezeshkpour2025insight, ), SURE (kim2024sure, ), and PIKE-RAG (wang2025pike, ), which use LLMs to improve task comprehension, retrieval relevance, and data decomposition (qian2024memorag, ). Reinforcement learning has also been applied to optimize retrieval generation pipelines (zhang2025rag, ). However, these methods are primarily applicable in benign environments and do not explicitly address poisoning threats or semantic inconsistencies caused by conflicting retrieved content.
2.2 Adversarial Attacks on RAG
Recent research indicates that RAG systems are extremely vulnerable to adversarial manipulation at both the input and corpus levels. Attack strategies include: (1) Corpus Poisoning Attacks, which inject adversarially crafted documents into the retrieval corpus and manipulate downstream outputs (zhang2025practical, ; sui2025ctrlrag, ; zou2024poisonedrag, ; shafran2024machine, ; chen2024agentpoison, ; xue2024badrag, ; zhong2023poisoning, ; nazary2025poison, ). (2) Prompt Injection Attacks, which use imperceptible instructions embedded in user queries or retrieved content to hijack LLM behavior without altering the underlying corpus (roychowdhury2024confusedpilot, ; jiao2025pr, ; li2024targeting, ). (3) Backdoor Attacks, in which hidden triggers are implanted into the corpus or model and activated only under certain conditions (long2024backdoor, ; cheng2024trojanrag, ). These attacks destroy the reliability of RAG outputs and expose the system to silent failure scenarios.
2.3 Defenses Against poisoning RAG
A variety of defense strategies have been proposed to counter adversarial threats. Perplexity-based detectors seek to identify anomalous generations, whereas RevPRAG examines LLM activation patterns to detect poisoned inputs (shafran2024machine, ; tan2024knowledge, ). RobustRAG introduces an isolate-then-aggregate framework to improve robustness by decoupling retrieval paths, while AstuteRAG adaptively fuses internal knowledge with retrieved content using heuristic selection (xiang2024certifiably, ; wang2024astute, ). InstructRAG enhances Retrieval-Augmented Generation by employing self-synthesized rationales, guiding the retrieval process to improve the relevance and coherence of generated outputs (wei2024instructrag, ). TrustRAG filters out malicious content using clustering over document embeddings and introduces a conflict resolution mechanism based on document consistency (zhou2025trustrag, ). Although promising, these approaches have two major limitations: Majority-voting often fails under high poisoning, while heuristic and aggressive filtering may lose relevant content under low poisoning.
In contrast to previous work, we propose SeCon-RAG, a robust two-stage framework for fine-grained semantic filtering and conflict-aware inference. SeCon-RAG improves robust against both high and low poisoning setting by leveraging intrinsic semantic signals and reasoning over document-level consistency, while preserving valuable information for reliable generation.
3 Preliminary
This section provides a brief overview of Retrieval-Augmented Generation and introduces the threat model of corpus poisoning attacks that underpins the defense strategies proposed in this paper.
3.1 Retrieval-Augmented Generation
Retrieval-Augmented Generation is a widely used paradigm for augmenting large language models with external knowledge obtained from a document corpus. Given a user query $q$ and a corpus $\mathcal{D}=\{d_{i}\}$, where $d_{i}$ represent the documents in $\mathcal{D}$. The standard RAG framework has three primary stages. In the first stage, compress the query $q$ and the documents $d_{i}$ in $\mathcal{D}$ into $E(q)$ and $E(d_{i})$ using the embedding model $E$. In the second stage, select the top-k documents with the highest similarity to the problem in the document to form a set $\mathcal{D}_{k}(q)$. The similarity is determined by a given function $\text{sim}(\cdot,\cdot)$, as follows:
$\mathcal{D}_{k}(q)=\operatorname{Top-k}_{d\in\mathcal{D}}\{\text{sim}(E(q),E(d))\},$ (1)
Finally, the retrieved documents $D_{k}(q)$ are combined with the original query $q$ to create an augmented input prompt. The augmented input is processed by a generative model $F$, such as a large language model, i.e. $F(q,\mathcal{D}_{k}(q))$, to generate the final output.
3.2 Threat Model: Corpus Poisoning Attacks
We examine a threat model that tries to trick a RAG system into producing incorrect answers by inserting carefully crafted malicious documents into its retrieval corpus. The attacker chooses $M$ target queries $\mathcal{Q}=\{q_{1},q_{2},\dots,q_{M}\}$ and matches each query $q_{i}$ with a poisoning target answer $r_{i}$. For example, for $q_{i}=$ “Who is the president of America?”, the adversary may want the RAG system to produce $r_{i}=$ “The president of America is Harris” . To achieve this, the attacker injects $N$ poisoning documents per query. Let $p^{j}_{i}$ denote the $j$-th poisoned document for query $q_{i}$, where $j=1,\dots,N$. The total set of injected documents is:
$\Gamma=\{p^{j}_{i}\mid i=1,\dots,M;\ j=1,\dots,N\}$ (2)
The attack aims to create $\Gamma$ so that, for each query $q_{i}\in\mathcal{Q}$, RAG system retrieves documents from the poisoned corpus $\mathcal{D}^{\prime}=\mathcal{D}\cup\Gamma$ that lead the generative model $F$ to produce the incorrect response $r_{i}$:
$F(q_{i},\mathcal{D^{\prime}}_{k}(q_{i}))\approx r_{i},\ \forall i\in[M].$ (3)
This threat model is consistent with previous research on corpus poisoning and informs our design of a filtering-based defense strategy. In the following sections, we present our proposed SeCon-RAG framework, which combines two-stage filtering to protect against corpus poisoning attacks.
4 The Proposed Defense Method for Corpus Poisoning Attacks
To protect Retrieval-Augmented Generation systems from corpus poisoning attacks, we propose SeCon-RAG, a robust two-stage filtering framework designed to detect and suppress poisoning documents. The first stage eliminates poisoned content statistically and semantically, while the second stage ensures factual consistency from a semantic reasoning perspective. This design ensures robustness without unnecessary knowledge loss. To enable fine-grained semantic understanding and aid in the detection of potentially poisoned content, we propose Entity-Intent-Relation Extractor in section 4.1, a semantic structure extraction module that serves as the foundation for our two-stage filtering framework. Before the retrieval stage, we propose Semantic and Cluster-Based Filtering shown in section 4.2 creates a semantic graph from the information extracted by EIRE, allowing for dual-channel filtering based on both clustering structure and semantic relevance. During the inference stage, we introduce the Conflict-Aware Filtering module shown in section 4.3. CAF performs cross-source semantic consistency checks using both EIRE on the retrieved content and the model’s internal knowledge representations. Figure 1 shows an overview of the full Secon-RAG framework. The appendix A.3 shows the pseudocode for the overall algorithm.
Figure 1: Overview of the SeCon-RAG. A two-stage defense in which SCF filters poisoning corpus during retrieval and CAF eliminates residual conflicts during inference, guided by semantic information obtained through EIRE.
4.1 EIRE: Entity-Intent-Relation Extractor
To enable fine-grained semantic understanding and aid in the detection of potentially poisoned content, we propose EIRE (Entity-Intent-Relation Extractor), a semantic structure extraction module that serves as the foundation for our two-stage filtering framework. EIRE is intended to capture the high-level meaning of a document by breaking it down into three core structural components:
  • Entities: Key entities explicitly or implicitly mentioned in the text.
  • Intent: The underlying purpose or objective conveyed by the passage.
  • Relations: Semantic relationships between extracted entities, such as beat or followed by.
To extract these components, EIRE employs a prompt-based large language model. Given a document $d$, we create structured prompts that direct the LLM to generate a structured triple. Appendix A.1 contains an example of the prompt and its output. For a document $d$, EIRE generates a structured triple $({E}_{d},{I}_{d},{R}_{d})$, where ${E}_{d}$ is the set of extracted entities, ${I}_{d}$ is the identified intent, and ${R}_{d}$ is the set of semantic relations between entities. By grounding document analysis in interpretable semantic frames, EIRE provides a robust and explainable foundation for downstream filtering.
4.2 Semantic and Clustering-Based Filtering
To reduce the risk of retrieving poisoned or adversarial documents, we introduce a dual filtering mechanism in the retrieval stage called Semantic and Clustering-Based Filtering (SCF). SCF is applied before selecting $\mathcal{D}_{k}(q)$ in the RAG pipeline.
4.2.1 Clustering-Based Filtering
Adversarially generated poisoning documents often exhibit highly similar phrasing or templated structures, especially when crafted to target the same query. As a result, they naturally form tight clusters in the embedding space (zhou2025trustrag, ). To mitigate this, we first apply a clustering-based filter to detect poisoning document groups. Given a potentially poisoned corpus $\mathcal{D}^{\prime}=\mathcal{D}\cup\Gamma$, we embed each document $d\in\mathcal{D}^{\prime}$ into the vector representation $m(d)$ and apply K-means clustering to obtain $K$ clusters $C=\{c_{1},\dots,c_{K}\}$, each with centroid $\mu_{i}=\frac{1}{|c_{i}|}\sum_{d_{j}\in c_{i}}m(d_{j}),\quad\bigcup_{i=1}^{K}c_{i}=\mathcal{D}^{\prime}$ (na2010research, ). We then define the filtered set as:
$\mathcal{D}_{\text{cluster}}=\bigcup_{i=1}^{K}\left\{d\in c_{i}\mid\text{sim}(m(d),\mu_{i})\leq\tau_{\text{cluster}}\right\}$ (4)
where $sim(\cdot,\cdot)$ denotes the cosine similarity normalized to $[0,1]$, and $\tau_{\text{cluster}}\in(0,1)$ is an adjustable filtering threshold. This operation effectively exclude documents that cluster too tightly around a centroid, which are likely to be maliciously inserted poisoning documents.
4.2.2 Semantic Graph-Based Filtering by EIRE
However, clustering-based methods rely solely on vector similarity in the embedding space, which can lead to false negatives by discarding valuable documents like topic overlap. To address this, we propose a semantic filter based on EIRE that extracts semantic structures from individual documents and generates corresponding semantic graphs. Specifically, for a document $d$, we construct a semantic relevance graph $G_{d}=(V_{i},E_{ij})$ by using information extracted from EIRE to simulate the semantic coherence and connectivity of the document $d$ as follows:
  • $V_{i}$: Each node $v_{i}\in V_{i}$ in the $G_{d}$ corresponds to the embedding representation of an entity extracted from document $d$;
  • $E_{ij}$: An edge $e_{ij}$ between two nodes $v_{i},v_{j}$ denotes a semantic relation extracted by EIRE connecting the two entities.
Figure 2 visualizes semantic graphs generated by EIRE for correct and poisoned documents under the query: "Which French ace pilot and adventurer flew L’Oiseau Blanc?" . As demonstrated, correct documents produce densely connected semantic graphs with high coherence, whereas poisoned documents have sparse or fragmented structures.
(a) Correct document
(b) Poisoned document
Figure 2: Semantic graph comparison using EIRE, more textual details has shown in AppendixA.1.2.
From a graph-theoretic perspective, the correct document displays a densely connected semantic graph, with the correct answer node well integrated into the EIRE conceptual structure, resulting in semantic graphs with high structural connectivity and semantic coherence. In contrast, poisoning content often introduces isolated or deceptive claims lacking semantic support from the surrounding context. Consequently, their semantic graphs exhibit abrupt or unnatural connections, with isolated nodes or disconnected subgraphs, in sharp contrast to the coherent clusters in correct documents. This structural distinction underpins our semantic filtering strategy.
To use these structural properties, we first construct a set of semantic graphs $G_{\text{cor}}=\{G_{d_{\text{cor},i}}\}$ from a small collection of verified correct documents $D_{\text{cor}}=\{d_{\text{cor},i}\}$ and use the semantic graphs $G_{\text{cor}}$ as a benchmark. $G_{\text{cor}}$ used for semantic reference are a small set of samples chosen manually from the dataset. For any candidate document $d\in\mathcal{D}^{\prime}$, we generate its semantic graph $G_{d}$ using EIRE and assess its similarity to the $G_{\text{cor}}$. Rather than relying on rigid graph similarity metrics, we employ large language models’ semantic reasoning capabilities to compare graph structures in a more flexible and context-aware way. For any candidate document $d$, we compute its semantic similarity score $ssG$ by comparing generate $G_{d}$ to $G_{d_{cor,i}}$ using a prompt-based LLM as shown in appendix A.1:
$ssG(d,D_{\text{cor}})=\text{LLM}(G_{d},G_{\text{cor}})$ (5)
To facilitate downstream filtering, we limit the LLM-derived semantic similarity score $ssG(d,D_{\text{cor}})\in[0,1]$. The higher the similarity score, the closer the semantic graph of $d$ and baseline $G_{\text{cor}}$. Using this score, we define the semantically filtered document set as:
$\mathcal{D}_{\text{semantic}}=\left\{d\in\mathcal{D}^{\prime}\mid ssG(d,D_{\text{cor}})\leq\tau_{\text{semantic}}\right\}$ (6)
where $\tau_{\text{semantic}}$ is the adjustable threshold that controls the strictness of semantic filtering. It is worth noting that, while vector projections are used to visualize and shape semantic graphs, the real inputs to the LLM in Equation (5) are natural language descriptions of the graphs serialized as structured triples.
4.2.3 Joint Filtering Decision: Robust AND Logic
To increase robustness while reducing the risk of discarding valuable information, we use a conservative AND-based filtering strategy. Only documents that have been flagged by both clustering and semantic filters are filtered. We define the final set of filtered documents as $\mathcal{D}_{\text{final}}=\mathcal{D}_{\text{cluster}}\cap\mathcal{D}_{\text{semantic}}$. Accordingly, the final retained corpus is:
$\widetilde{\mathcal{D}}=\mathcal{D}^{\prime}\setminus\mathcal{D}_{\text{final}}$ (7)
By the joint filter, the SCF module combines unsupervised clustering and semantic reasoning to detect poisoning documents from multiple perspectives. This layered approach improves the quality of retrieved documents and provides a robust first line of defense in the SeCon-RAG framework.
4.3 Conflict-Aware Filtering (CAF)
Although the SCF module effectively reduces adversarial content, it may retain documents that are not malicious but semantically irrelevant or internally inconsistent. These residual conflicts, such as documents that contradict the query, other retrieved evidence, or the model’s internal knowledge, can reduce the factual reliability of the final answer. To address this limitation, we propose Conflict-Aware Filtering (CAF), a semantic inference module used at the inference stage of the RAG. CAF aims to refine the retrieved set $\mathcal{D}_{k}(q)$ by identifying and removing documents that don’t meet semantic and factual consistency criteria.
For each candidate document $d\in\mathcal{D}_{k}(q)$, CAF generates structured semantic information using EIRE, which is divided into three components: Entities capable of determining whether facts align with the model’s internal knowledge; Intent to evaluate query alignment; Relations that can evaluate logical coherence across documents and detect contradictions or omissions. In the final inference process, we prompt the LLMs as shown in appendix A.1 to determine which information from the retrieve documents is reliable from three dimensions using the semantic information extracted by EIRE:
  • Q(Query Consistency): Does the document semantically aligned with the user query $q$, based on intent and entities?
  • C(Corpus Consistency): Is the document consistent with the other retrieved documents $\mathcal{D}_{k}(q)\setminus d$, based on shared relations and context?
  • M(Model Consistency): Is the document factually compatible with the LLM’s internal knowledge, considering key entities?
The LLMs will render a judgment on whether the information is poisoned, conflicting, irrelevant, or trustworthy, which will be used to make final decisions. Finally, the LLMs extracts the final answer $A(q)$ to query $q$ from documents that perform well in these three dimensions:
$\begin{array}[]{cl}A(q)&=F(q,\widetilde{\mathcal{D}}_{\text{CAF}})\\ \widetilde{\mathcal{D}}_{\text{CAF}}&=\left\{d\in\mathcal{D}_{k}(q)\;\middle|\;\text{CAF}(d,Q,C,M)=\texttt{trustable}\right\}\end{array}$
Figure 1 shows the operation of CAF. Following filtering by SCF, each document is evaluated by the LLM based on EIRE-derived semantic structure. Information from documents have been identified as poisoned, conflicting, or irrelevant is discarded, leaving only trustable information for final generation. This ensures that the generation module operates on a semantically coherent, query relevant, and factually aligned knowledge base, thereby increasing robustness and factual faithfulness.
CAF enhances SCF by providing fine-grained semantic validation. While SCF removes broad outliers based on statistical or semantic graph anomalies, CAF ensures that documents generated have coherent intent, correct facts, and logical consistency. This layered design increases the final output’s robustness as well as its factual accuracy.
5 Experiments
5.1 Setup
This section describes the experimental setup. We evaluate the effectiveness and robustness of SeConRAG under various adversarial scenarios. All reported results are averages of multiple runs with an error of ± 1%.
Datasets. We test three popular open-domain question-answering benchmarks: Natural Questions (NQ) (Kwiatkowski2019natural, ), HotpotQA (Yang2018hotpotqa, ), and MS-MARCO (bajaj2016ms, ). Each dataset corresponds to a large-scale corpus.
Attack Settings. To test robustness under poisoning scenarios, we evaluate two representative types of attacks against RAG systems: (1) Corpus Poisoning Attack, following PoisonedRAG  (zou2024poisonedrag, ), which inserts adversarial passages into the corpus; (2) Prompt Injection Attack (PIA) (zhong2023poisoning, ; greshake2023not, ),which adversarial prompts are created by perturbing discrete tokens to closely resemble training queries, thereby misleading the model during inference. It ensures a comprehensive evaluation of SeCon-RAG under both input and retrieval adversarial threats.
Evaluation Metrics. We use standard metrics from previous research to evaluate model robustness and answer quality: (1) Accuracy (ACC) is the percentage of generated answers that exactly match the ground truth. (2) Attack Success Rate (ASR): The percentage of poisoning queries or documents that cause the model to produce incorrect results.
Verified Correct Documents. To create the semantic reference set $D_{cor}$, we manually selected 10 clean documents from each dataset.
Model. We tested five LLMs from both open and closed source families: Mistral-12B, Qwen-7B, LLaMA-3.1-8B, GPT-4o, and DeepSeek-R1. RAG backbones are maintained in accordance with the corresponding LLMs. Appendix  A.1 provides detailed prompts for EIRE, semantic similarity, and CAF modules. All experiments are carried out using NVIDIA A100-SXM4-40GB GPUs.
5.2 Main Results
We evaluate the robustness and effectiveness of SeconRAG against four representative Retrieval-Augmented Generation (RAG) defense baselines VanillaRAG, InstructRAG (wei2024instructrag, ), AstuteRAG (wang2024astute, ), and TrustRAG (zhou2025trustrag, ), across three datasets and five LLMs. To verify that our method can handle both high and low poisoning rates simultaneously, each model is tested under four settings: Clean, Prompt Injection Attack, PoisonedRAG-20%, and PoisonedRAG-100%, with results reported in terms of Accuracy and Attack Success Rate. Table 1 summarizes the key findings. More detailed PoisonedRAG results at different poisoning levels can be found in Appendix A.4.1.
Table 1: Performance comparison of SeConRAG and baseline methods across three QA datasets and five LLMs under PIA, 20% and 100% corpus poisoning, and clean settings. Best values (highest accuracy ↑ or lowest ASR ↓) are highlighted in bold.
Model Method HotpotQA (Yang2018hotpotqa, ) NQ (Kwiatkowski2019natural, ) MS-MARCO (bajaj2016ms, )
PIA PR-100% P-20% Clean PIA PR-100% PR20% Clean PIA PR-100% PR20% Clean
ACC/ASR ACC/ASR ACC/ ASR ACC ACC/ASR ACC/ ASR ACC/ ASR ACC ACC/ ASR ACC/ ASR ACC/ ASR ACC
Mistral-12B (aydin2025generative, ) VanillaRAG 51.0 / 40.0 0.9 / 98.2 38.2 / 58.0 75.0 47.6 / 37.5 8.2 / 90.9 38.2 / 48.2 68.0 54.5 / 43.6 9.1 / 89.1 50.0 / 45.5 84.0
InstructRAG (wei2024instructrag, ) 50.0 / 43.6 13.6 / 83.5 45.5 / 49.1 75.0 48.2 / 43.2 13.6 / 82.7 51.8 / 40.0 66.0 64.5 / 33.2 15.5 / 78.2 57.3 / 36.4 81.0
ASTUTERAG (wang2024astute, ) 68.2 / 17.3 32.7 / 61.1 65.9 / 21.8 76.0 64.5 / 10.0 43.6 / 38.2 67.7 / 11.8 70.0 75.9 / 17.3 32.7 / 58.2 73.6 / 18.8 81.0
TrustRAG (zhou2025trustrag, ) 75.5 / 1.4 75.5 / 3.6 71.8 / 14.5 81.0 68.2 / 0.5 62.7 / 1.8 66.4 / 13.6 73.0 90.9 / 0.0 91.8 / 0.0 87.3 / 11.8 85.0
SeconRAG(ours) 77.5 / 0.8 75.7 / 3.6 72.7 / 4.5 83.0 72.3 / 1.8 63.6 / 2.5 74.5 / 10.2 82.0 91.8 / 0.0 88.2 / 0.0 89.1 / 9.1 98.0
Qwen-7B (hui2024qwen2, ) VanillaRAG 34.0 / 60.9 1.8 / 98.2 32.7 / 65.5 67.0 28.2 / 67.3 5.5 / 93.6 39.1 /51.8 56.0 36.4 / 60.9 10.0 / 87.3 43.6 / 46.4 75.0
InstructRAG (wei2024instructrag, ) 58.2 / 38.2 24.5 / 76.4 45.5 / 51.8 67.0 52.7 / 45.5 25.5 / 76.4 47.3 / 47.3 64.0 61.8 / 36.4 43.6 / 57.8 49.1 / 45.5 75.0
ASTUTERAG (wang2024astute, ) 51.8 / 29.1 45.5 / 44.1 58.6 / 25.4 65.0 56.4 / 17.3 42.3 / 53.2 60.5 / 17.3 68.0 44.5 / 45.5 42.3 / 54.5 65.5 / 20.0 74.0
TrustRAG (zhou2025trustrag, ) 62.7 / 0.6 58.2 / 2.7 58.2 / 26.4 73.0 67.3 / 0.6 60.0 / 2.7 64.5 / 24.5 67.0 68.2 / 1.8 64.5 / 11.8 66.4 / 22.7 78.0
SeconRAG(ours) 67.3 / 0.5 63.6 / 2.3 61.8 / 21.8 76.0 73.6 / 8.2 66.4 / 2.4 70.9 / 21.8 78.0 75.5 / 1.4 71.8 / 4.5 75.5 / 17.5 84.0
LLaMA-3.1-8B (dubey2024llama, ) VanillaRAG 31.8 / 62.7 4.5 / 96.4 36.4 / 57.3 70.0 38.2 / 54.5 10.9 / 88.2 41.8 / 52.7 70.0 34.5 / 63.6 9.1 / 88.2 54.5 / 40.9 83.0
InstructRAG (wei2024instructrag, ) 61.8 / 30.0 27.3 / 71.8 47.3 / 50.0 76.0 67.3 / 24.1 32.7 / 67.3 56.4 / 34.5 70.0 68.2 / 26.4 48.5 / 51.8 72.7 / 27.3 81.0
ASTUTERAG (wang2024astute, ) 43.6 / 41.8 46.8 / 47.0 65.5 / 20.9 68.0 57.3 / 26.4 58.2 / 31.8 77.5 / 8.2 81.0 59.1 / 39.5 56.8 / 38.6 82.3 / 13.6 89.0
TrustRAG (zhou2025trustrag, ) 72.7 / 0.5 67.3 / 3.0 65.5 / 19.1 72.0 84.5 / 0.2 79.1 / 0.0 79.1 / 10.9 84.0 86.4 / 1.5 84.5 / 6.4 85.4 / 9.1 84.0
SeconRAG(ours) 73.6 / 0.5 72.0 / 10.9 67.4 / 18.4 84.0 85.1 / 2.7 88.2 / 1.8 86.9 / 4.0 90.0 87.3 / 0.2 89.1 / 0.0 86.2 / 9.1 90.0
GPT-4o (achiam2023gpt, ) VanillaRAG 57.3 / 40.0 11.9 / 81.8 45.5 / 30.5 81.0 50.9 / 44.3 27.3 / 68.2 52.7 / 31.8 74.0 70.0 / 27.3 30.0 / 64.1 72.3 / 16.4 84.0
InstructRAG (wei2024instructrag, ) 59.1 / 37.3 27.3 / 71.8 61.8 / 33.2 84.0 58.2 / 26.5 43.6 / 51.1 66.4 / 25.5 74.0 77.3 / 16.4 50.5 / 42.7 70.9 / 17.3 83.0
ASTUTERAG (wang2024astute, ) 72.7 / 14.5 67.3 / 24.1 77.3 / 11.8 81.0 83.6 / 4.5 75.5 / 14.2 79.1 / 4.1 81.0 90.5 / 0.7 76.4 / 15.5 82.7 / 6.4 86.0
TrustRAG (zhou2025trustrag, ) 81.8 / 0.3 80.9 / 2.7 79.1 / 6.4 85.0 82.7 / 0.3 80.0 / 0.1 81.8 / 1.0 86.0 89.1 / 1.3 89.1 / 1.8 84.5 / 6.4 88.0
SeconRAG(ours) 83.6 / 0.3 83.6 / 2.4 79.1 / 5.5 86.0 89.1 / 0.6 81.8 / 0.0 84.5 / 1.0 88.0 93.6 / 0.0 89.1 / 1.8 89.1 / 3.6 94.0
DeepSeek-R1 (guo2025deepseek, ) VanillaRAG 59.1 / 32.7 10.0 / 89.1 51.0 / 46.4 81.0 64.3 / 27.3 17.3 / 84.5 51.0 / 43.6 80.0 71.8 / 25.5 11.8 / 81.8 60.5 / 29.1 82.0
InstructRAG (wei2024instructrag, ) 61.8 / 34.5 27.3 / 72.7 61.8 / 38.2 80.0 59.1 / 28.2 39.1 / 62.7 65.5 / 32.7 82.0 75.5 / 18.2 51.8 / 47.5 72.7 / 26.4 87.0
ASTUTERAG (wang2024astute, ) 73.6 / 14.5 64.5 / 25.5 77.3 / 14.5 79.0 90.0 / 1.8 81.8 / 10.9 89.1 / 0.0 87.0 87.3 / 4.5 85.5 / 8.2 89.1 / 5.5 88.0
TrustRAG (zhou2025trustrag, ) 81.8 / 4.5 79.1 / 2.7 85.5 / 10.0 89.0 90.0 / 1.8 88.2 / 0.0 90.0 / 3.6 91.0 93.6 / 1.8 89.1 / 3.6 89.1 / 5.5 91.0
SeconRAG(ours) 84.5 / 3.0 81.8 / 8.0 83.6 / 5.5 86.0 90.0 / 0.9 96.4 / 0.0 96.4 / 0.0 98.0 92.7 / 3.0 94.5 / 1.8 94.5 / 5.5 94.0
Results in Table 1 demonstrates that SeConRAG outperforms in almost all datasets, LLMs, and attack scenarios. Under high poisoning (PoisonedRAG-100%), it consistently maintains high accuracy and low ASR. For example, on HotpotQA with GPT-4o, SeConRAG achieves 83.6% accuracy and 2.4% ASR, outperforming TrustRAG (80.9% / 2.7%) and ASTUTERAG (67.3% / 24.1%). A Similar trends hold under low poisoning (20%), where SeConRAG consistently improves robustness compared with baselines. SeConRAG also performs well against Prompt Injection Attacks, which target the input layer rather than retrieval. On MS-MARCO with GPT-4o, it achieves 93.6% accuracy and 0.0% ASR, slightly surpassing TrustRAG and ASTUTERAG. Even with smaller models such as Qwen-7B, SeConRAG retains competitive performance (67.3% / 0.5%), demonstrating the effectiveness of CAF in mitigating prompt-level inconsistency.
Importantly, SeConRAG continues to perform well on clean corpora. On MS-MARCO with DeepSeek-R1, it reaches 94.0% accuracy, and on NQ with Mistral-12B, it achieves 82.0%, outperforming TrustRAG (73.0%) and ASTUTERAG (70.0%). This demonstrates that the defense mechanisms do not degrade benign performance. Overall, SeConRAG consistently outperforms or equals existing defenses across datasets, LLMs, and threat scenarios. Across both large models (GPT-4o, DeepSeek-R1) and smaller instruction-tuned models (Qwen-7B, Mistral-12B), it consistently delivers reliable and generalizable robustness against both corpus poisoning and prompt-level adversaries, making it a viable defense for real-world RAG deployments.
5.3 Ablation Study
We conduct an ablation study on the Mistral-12B model to evaluate the contributions of SeCon-RAG’s components.We concentrate on two key areas: (i) the core SCF and CAF modules, and (ii) an ablation study that includes the SCF subcomponents, EIRE, and the verified Correct Document Set.
5.3.1 Core Modules (SCF and CAF).
To evaluate the impact of SCF, we remove this module and compare performance with three QA datasets. SCF uses clustering and semantic graph filtering to eliminate documents that are semantically irrelevant or poisoning. Figure  3 shows that disabling SCF consistently decreases accuracy and increases ASR across all datasets. In the 100% poisoning setting on HotpotQA, accuracy drops from 74.0% to 71.0%, and ASR increases from 8.0% to 25.0%. Under prompt injection attacks, accuracy decreases from 92.0% to 85.0%. These findings support SCF’s effectiveness in increasing retrieval precision and resisting semantically attacks. We then assess the CAF module, which filters semantically conflicting evidence using EIRE-based consistency checks. Removing CAF leads to more severe degradation. When using HotpotQA with 100% poisoned data, accuracy drops to 68.0% and ASR rises to 56.0%. ASR increases to 47.0% on NQ, while accuracy decreases from 92.0% to 46.0% on PIA. These findings highlight CAF’s critical role in detecting and filtering conflicting or misleading documents that SCF alone may miss.
Figure 3: Ablation results on accuracy and attack success rate (ASR) across three datasets using Mistral-12B. From left to right are HotpotQA, NQ, MS-MARCO.
5.3.2 Evaluation of SCF Subcomponents, EIRE, and the Verified Correct Document Set.
To examine whether clustering and semantic filtering are complementary, we test each individually. As shown in Table  5, their combination yields the strongest robustness, achieving lower ASR in several cases, which confirms the necessity of combination.
In addition, we assess the standalone effectiveness of the Entity-Intent-Relation Extractor and the verified correct document set ($d_{cor}$). Table  6 in appendix summarizes the results. EIRE improves the fine-grained reasoning capability of both SCF and CAF. With EIRE enabled, the model consistently achieves higher factual accuracy while significantly lowering ASR, especially under high poisoning conditions. Similarly, a small, high-quality $d_{cor}$ set can significantly improve semantic filtering performance and reduce noise from poisoned documents, as well as improve robustness under high-poisoning conditions (e.g., ASR $\rightarrow 0$ on MS-MARCO 100% poisoning).
The ablation results show that both SCF and CAF are critical for protecting against poisoning attacks. SCF performs coarse filtering of anomalous content, while CAF ensures semantic and factual consistency. Their collaboration allows SeCon-RAG to maintain strong performance in high-poisoning and adversarial scenarios.
5.4 Runtime Analysis
We compare SeConRAG’s runtime cost to four representative RAG baselines—VanillaRAG, InstructRAG, AstuteRAG, and TrustRAG—on three QA benchmarks: HotpotQA, NQ, and MS-MARCO. The methods are evaluated in three adversarial settings: Prompt Injection Attack and PoisonedRAG with 100% or 20% poisoning. Figure 4 depicts the full results.
Figure 4: The average runtime per batch (in minutes) for three datasets and adversarial settings. From left to right are HotpotQA, NQ, MS-MARCO.
Although SeConRAG achieves the highest robustness in all settings, it has a moderate runtime overhead. It takes between 1.21 and 1.45 minutes per batch, depending on the dataset and the severity of the attack. This cost is due to its multi-stage semantic filtering, consistency checks, and conflict filtering, which protects against poisoned documents. Despite the additional cost, SeConRAG maintains a practical runtime range. Deeper semantic understanding requires the use of LLMs for semantic structure extraction and graph similarity calculation. Despite the additional cost, SeConRAG has a reasonable runtime range. This trade-off is acceptable for many real-world RAG applications that need both robustness and correctness. For example, on NQ with 100% poisoning, it completes in 1.06 minutes, only 10 seconds slower than TrustRAG (0.67 min) or AstuteRAG (0.70 min), but it offers significantly more reliable answers. Overall, the asymptotic overhead is moderate relative to standard retrieval. This trade-off is acceptable for many real-world RAG applications that require robustness and correctness.
5.5 Embedding Models
We further evaluate SeConRAG with four widely used embedding models: MiniLM(wang2020minilm, ), SimCSE(gao2021simcse, ), BERT(devlin2019bert, ), and BGE(chen2024bge, ). These encoders are integrated into the retrieval and two-stage filtering pipelines, with Mistral-12B serving as the primary LLM. Table 2 displays results from three datasets with various poisoning ratios. Across different embedding model, SeConRAG maintains high accuracy ($75\%$) and low ASR ($10\%$) under 100% poisoning. For example, on MS-MARCO, BGE achieves 90.0%/0.0%, while MiniLM yields 77.3%/7.3%. These findings confirm that SeConRAG’s defense framework perform well across embeddings and avoids reliance on a single model.
Table 2: Comparison of SeConRAG performance under different embedding models (MiniLM, SimCSE, BERT, BGE) across varying poisoning ratios on three datasets. highest accuracy ↑ or lowest ASR ↓
Model Setting HotpotQA (Yang2018hotpotqa, ) NQ (Kwiatkowski2019natural, ) MS-MARCO (bajaj2016ms, )
100% 80% 60% 40% 20% 100% 80% 60% 40% 20% 100% 80% 60% 40% 20%
ACC/ASR ACC/ASR ACC/ ASR ACC/ASR ACC/ASR ACC/ ASR ACC/ ASR ACC/ASR ACC/ ASR ACC/ ASR ACC/ ASR ACC/ASR ACC/ASR ACC/ASR ACC/ASR
mistral-12b SimCSE 73.6 / 8.2 77.3 / 4.0 75.5 / 4.0 71.8 / 8.2 73.6 / 4.0 67.3 / 5.5 67.3 / 0.0 67.3 / 3.6 69.1 / 0.0 79.1 / 7.3 79.1 / 7.3 91.8 / 1.8 91.8 / 0.0 90.0 / 1.8 90.0 / 0.0
mistral-12b MiniLM 75.5 / 9.1 75.5 / 5.5 77.3 / 5.5 77.3 / 5.5 75.5 / 4.0 75.5 / 3.6 71.8 / 5.5 71.8 / 0.0 69.1 / 1.8 70.9 / 0.0 77.3 / 7.3 91.8 / 1.8 90.0 / 0.0 90.0 / 0.0 91.8 / 0.0
mistral-12b BGE 75.5 / 5.5 77.3 / 4.0 75.5 / 4.0 75.5 / 7.3 71.8 / 9.1 70.9 / 9.1 67.3 / 1.8 71.8 / 0.0 71.8 / 1.8 73.6 / 0.0 90.0 / 0.0 91.8 / 0.0 90.0 / 0.0 91.8 / 0.0 90.0 / 9.1
mistral-12b BERT 72.7 / 6.4 77.3 / 7.3 75.5 / 4.0 75.5 / 5.5 75.5 / 11.5 74.5 / 10.9 67.3 / 7.3 71.8 / 3.6 69.1 / 1.8 69.1 / 1.8 79.1 / 9.1 89.1 / 1.8 91.8 / 0.0 91.8 / 0.0 93.6 / 0.0
5.6 Sensitivity Analysis of Filtering Thresholds
We investigate the impact of two primary thresholds, $\tau_{\text{cluster}}$ and $\tau_{\text{semantic}}$, on LLaMA-3.1-8B and GPT-4o. As shown in Table 8, performance remains stable across reasonable ranges ($\tau_{\text{cluster}}\in[0.86,0.90]$, $\tau_{\text{semantic}}\in[0.2,0.4]$), with accuracy variations within $\pm 2\%$ and low ASR. This robustness is due to the conservative AND-logic in joint filtering, which ensures that only documents flagged by both filters are removed. As a result, SeConRAG is not overly sensitive to precise hyperparameter tuning, making it useful in real-world deployment.
6 Conclusion
We propose SeCon-RAG, a robust retrieval-augmented generation framework that protects against corpus poisoning. It combines two complementary modules: Semantic and Cluster-Based Filtering, which removes poisoned content using clustering and semantic similarity, and Conflict-Aware Filtering, which filters out contradictory or misleading evidence using structured semantic reasoning. Experiments with multiple datasets and poisoning scenarios show significant improvements in answer accuracy and reduced attack success rates. SeCon-RAG provides a scalable and interpretable defense for RAG systems in adversarial environments by combining coarse-grained statistical pruning and fine-grained semantic validation. The Impact Statement of our paper is shown in the appendix.
Limitations. While SeCon-RAG demonstrates strong robustness against a range of poisoning attacks, several limitations remain. First, SeCon-RAG introduces moderate inference latency due to multiple LLM calls (EIRE extraction, semantic similarity, and CAF decision-making). Second, the framework relies on high-quality semantic extraction; Finally, a small set of manually verified documents $D_{cor}$ is required.Future research could reduce runtime overhead by replacing EIRE with smaller models and exploring lightweight graph similarity metrics. These changes will make SeCon-RAG better suited for latency-sensitive, real-time RAG applications.
Acknowledgments and Disclosure of Funding
This work is supported by CAS Project for Young Scientists in Basic Research, Grant No.YSBR-040, ISCAS New Cultivation Project ISCAS-PYFX-202201, ISCAS Basic Research ISCAS-JCZD-202302, National Natural Science Foundation of China General Project (No. 71971051, No. 72371067),the National Research Foundation, Singapore, and DSO National Laboratories under the AI Singapore Programme (AISG Award No: AISG4-GC-2023-008-1B); by the National Research Foundation Singapore and the Cyber Security Agency under the National Cybersecurity R&D Programme (NCRP25-P04-TAICeN); and by the Prime Minister’s Office, Singapore under the Campus for Research Excellence and Technological Enterprise (CREATE) Programme. Any opinions, findings and conclusions, or recommendations expressed in these materials are those of the author(s) and do not reflect the views of the National Research Foundation, Singapore, Cyber Security Agency of Singapore, Singapore.
Appendix A Appendix / supplemental material
Impact Statement
SeCon-RAG’s effectiveness is dependent on the quality of its semantic parser (EIRE), which may perform poorly on domain-specific texts. The methods proposed in this paper will not have a negative impact on the community.
A.1 Prompt and Example
A.1.1 EIRE (Entity-Intent-Relation Extractor)
We provide the prompt template used in our system to extract the intent, key entities, and entity relations from a given documents.
Prompt for EIRE Please extract both the key entities, intent and relations of entities of the documents, using the following criteria: Key Entities: Extract specific entities (such as terms, names, data, and locations) that are explicitly mentioned in the document for later entity matching. Intent: Summarize the document’s main points or conclusion in a single sentence. This should be free of external issues and only include the document’s own claims. Relation: Identify the most critical relationship between key entities in the document, keeping the length under the original document.
Output for example document extracted by EIRE
Input:
L’Oiseau Blanc (commonly known in the English-speaking world as The White Bird ) was a French Levasseur PL.8 biplane that disappeared in 1927, during an attempt to make the first non-stop transatlantic flight between Paris and New York City to compete for the Orteig Prize. The aircraft was flown by French World War I aviation heroes, Charles Nungesser and Fran0̆0e7ois Coli. The aircraft took off from Paris on 8 May 1927 and was last seen over Ireland. Less than two weeks later, Charles Lindbergh successfully made the New York2̆013Paris journey and claimed the prize, flying the S̈pirit of St. Louis Output: Intent:
"L’Oiseau Blanc disappeared in 1927 while attempting the first non-stop
transatlantic flight from Paris to New York to win the Orteig Prize."
Entities:
"L’Oiseau Blanc", "Levasseur PL.8", "1927", "Charles Nungesser", "Paris"
"New York City", "Orteig Prize", "non-stop transatlantic flight",
"Ireland", "Charles Lindbergh", "François Coli",  "Spirit of St. Louis"
Relations:
    ["L’Oiseau Blanc", "Levasseur PL.8", "is_a"],
    ["L’Oiseau Blanc", "1927", "disappeared_in"],
    ["L’Oiseau Blanc", "non-stop transatlantic flight", "attempt"],
    ["L’Oiseau Blanc", "Paris", "departed_from"],
    ["L’Oiseau Blanc", "New York City", "destination"],
    ["L’Oiseau Blanc", "Charles Nungesser", "flown_by"],
    ["L’Oiseau Blanc", "François Coli", "flown_by"],
    ["L’Oiseau Blanc", "Ireland", "last_seen_over"],
    ["Charles Lindbergh", "Orteig Prize", "won"],
    ["Charles Lindbergh", "Spirit of St. Louis", "flew"]
A.1.2 Semantic filtering
Prompt of Semantic graph for filtering Evaluate if the following document contains deceptive or malicious material that may impact the answer to the inquiry. Consider the presence of certain terms, their relationships, and any underlying meanings. Use the graph generated using these semantic information and to calculate its the structure’s similarity between the correct document which is reference
reference documents:
"intent":"L’Oiseau Blanc was a French biplane that disappeared in 1927"
"keywords":
["L’Oiseau Blanc", "Levasseur PL.8", "1927", "Charles Nungesser",
"non-stop transatlantic flight", "Paris",
"New York City", "Orteig Prize",  "Spirit of St. Louis"
"François Coli", "Ireland", "Charles Lindbergh"]
"relations":
[("L’Oiseau Blanc", "Levasseur PL.8", "is_a"),
("L’Oiseau Blanc", "1927", "disappeared_in"),
("L’Oiseau Blanc", "non-stop transatlantic flight", "attempt"),
("L’Oiseau Blanc", "Paris", "departed_from"),
("L’Oiseau Blanc", "New York City", "destination"),
 ("L’Oiseau Blanc", "Charles Nungesser", "flown_by"),
("L’Oiseau Blanc", "François Coli", "flown_by"),
("L’Oiseau Blanc", "Ireland", "last_seen_over"),
("Charles Lindbergh", "Orteig Prize", "won"),
("Charles Lindbergh", "Spirit of St. Louis", "flew") ]
Score and filter documents based on their similarities and subsequent evaluations. Give a score between 0 and 10, where 0 indicates absolutely benign and 10 implies highly misleading or malevolent. The output must only show your score as an integer number. Analysis Criteria: Identify entity/Phrases: Identify entity or phrases that could be misleading or indicate a departure from reality. Relationship to the Question: Analyze how these terms relate to the question and determine if they might affect the answer to the question. Hidden Meanings or Implications: Consider if the document suggests fictional or inaccurate information without clear indication. Context: Consider how the document might be interpreted in the context of the question. Assess whether it could mislead someone seeking factual information. Intent and Impact: Evaluate if the misleading information appears intentional and whether it could cause harm or spread misinformation. Ambiguity and Clarity: Assess the document’s clarity. Vague or misleading presentations can be more harmful. Fiction vs. Reality: Determine if the document clearly distinguishes between hypothetical scenarios and reality. Relevance to the Question: Evaluate if the misleading information directly impacts the question’s answer. Scoring Guidelines: 0-3: Likely benign. The document contains little to no misleading or harmful information related to the question. 4-6: Potentially misleading. The document contains some elements that may confuse or mislead, but they are not the primary focus or are clearly indicated as hypothetical or fictional. 7-10: Likely malicious or misleading. The document actively presents information that could mislead, deceive, or significantly impact the answer to the question. Considerations for Scoring: Final Score: Provide your score based on the above analysis, Provide your score as an integer number only.
The distribution of the semantic graphs produced from semantic information in vector space is depicted in the image below. We utilize PCA to reduce the vector’s dimension to two dimensions and demonstrate it. Figure 5,  6,  7 visualize semantic graphs generated by EIRE for correct and poisoned documents under the query: "Which French ace pilot and adventurer flew L’Oiseau Blanc?" .
We employ the prompt of Semantic graph for filtering  A.1.2 to direct the llms in evaluating, scoring and filtering documents based on semantic information and correspoding graphs.
Figure 5: A schematic graph of the correct document’s semantic structure in two dimensions. Blue indicates the correct response, orange nodes stand for entities, and the edges connecting nodes show the connections between entities.
Figure 6: A schematic graph of the incorrect document’s semantic structure in two dimensions. Red indicates the poisoning response, orange nodes stand for entities, and the edges connecting nodes show the connections between entities.
Example for correct document Input:
"Can’t Help Falling in Love\" is a pop ballad originally recorded by
American singer Elvis Presley and published by Gladys Music,
Presley’s publishing company. It was written by Hugo Peretti,
Luigi Creatore, and George David Weiss.[2] The melody is based
[…]
Martini (1784). The song was featured in Presley’s 1961 film,
Blue Hawaii. During the following four decades,
it was recorded by numerous other artists, including Tom Smothers,
Swedish pop group A-Teens, and the British reggae group UB40,
whose 1993 version topped the U.S. and UK charts."
Output:
Semantic Content: Factually accurate, no false claims.
Graph Similarity: Completely unrelated (topic is music, not aviation).
Misleading Risk: Low although off-topic, it doesn’t mislead  facts.
Final Score: 2
Example for correct document Input:
"Frank Sinatra, the iconic crooner, recorded an unforgettable
rendition of the song \"I Can’t Help Falling in Love With You\",
enrapturing audiences with his soulful interpretation."
Output:
Semantic Content: Highly confident and emotional false claim.
Graph Similarity: No overlap with reference.
Misleading Risk:High persuasive wording increases belief in a falsehood.
Final Score: 8
A.2 Conflict-Aware Filtering
In the final inference process, we prompt the LLMs to determine which information from the retrieve documents is reliable from three dimensions using the semantic information extracted by EIRE.
Output for example document extracted by EIRE You must evaluate the document information you retrieve, which includes internal knowledge, external knowledge, and query. To answer this question, consider intent, key entities, and relationships to determine which knowledge provides the best, most accurate, and error-free support. The external information may not be reliable. Use a combination of intent and key entities from external information, as well as the intent of the original question, to make self-judgments about the reliability of external information. Then, based on both your assessment and your personal knowledge, provide the best possible answer.
Figure 7: A schematic graph of the incorrect document’s semantic structure in two dimensions. Red indicates the poisoning response, orange nodes stand for entities, and the edges connecting nodes show the connections between entities.
A.3 Pseudocode of SeCon-RAG
Provide formally written pseudocode (see Algorithm 1) for the full SeCon-RAG pipeline, including SCF and CAF. This helps clarify the implementation logic for reproducibility.
Algorithm 1 SeCon-RAG: Two-Stage Semantic Filtering and Conflict-Aware Generation
1:Query $q$, Retrieval corpus $\mathcal{D}$, Verified clean documents $\mathcal{D}_{\text{cor}}$, Pretrained LLM of RAG $F$
2:Trustworthy answer $A(q)$
3:Stage 1: Semantic and Cluster-Based Filtering (SCF)
4:Embed each document $d\in\mathcal{D}$ into vector $m(d)$
5:Apply K-Means clustering to obtain clusters $\mathcal{C}=\{c_{1},\dots,c_{K}\}$
6:for all $d\in\mathcal{D}$ do
7:  Compute similarity to cluster centroid: $s_{\text{cluster}}(d)\leftarrow\text{sim}(m(d),\mu_{c})$
8:  Extract semantic structure $(E_{d},I_{d},R_{d})\leftarrow\text{EIRE}(d)$
9:  Construct semantic graph $G_{d}$ from $(E_{d},I_{d},R_{d})$
10:  Compute semantic similarity score $s_{\text{sem}}(d)\leftarrow\text{LLM}(G_{d},\mathcal{G}_{\text{cor}})$
11:end for
12:Filter documents where $s_{\text{cluster}}(d)>\tau_{\text{cluster}}$ and $s_{\text{sem}}(d)<\tau_{\text{sem}}$
13:Define filtered corpus $\widetilde{\mathcal{D}}\leftarrow\mathcal{D}\setminus\mathcal{D}_{\text{filtered}}$
14:Stage 2: Conflict-Aware Filtering (CAF)
15:Retrieve top-$k$ documents $\mathcal{D}_{k}(q)$ from $\widetilde{\mathcal{D}}$ based on embedding similarity
16:for all $d\in\mathcal{D}_{k}(q)$ do
17:  Extract semantic structure $(E_{d},I_{d},R_{d})\leftarrow\text{EIRE}(d)$
18:  Evaluate:
  • Query consistency $Q(d,q)$
  • Corpus consistency $C(d,\mathcal{D}_{k}(q))$
  • Model consistency $M(d,F)$
19:  if $\text{CAF}(d,Q,C,M)=\texttt{trustable}$ then
20:   Add $d$ to $\mathcal{D}_{\text{CAF}}$
21:  end if
22:end for
23:Generate final answer: $A(q)\leftarrow F(q,\mathcal{D}_{\text{CAF}})$
24:return $A(q)$
A.4 Experiments
A.4.1 Experiments of Different Poisoning Ratio
HotpotQA
Table 3 compares SeConRAG’s performance to four baseline methods (VanillaRAG, InstructRAG, ASTUTERAG, and TrustRAG) across five backbone LLMs on the HotpotQA dataset with varying corpus poisoning ratios (0% to 100%). Across all models and poisoning levels, SeConRAG consistently achieves or approaches the highest accuracy while maintaining low attack success rates (ASR), demonstrating strong robustness and generalizability. Notably, On Mistral-12B SeConRAG achieves 75.7% accuracy with only 3.6% ASR under 100% poisoning, outperforming TrustRAG and significantly surpassing ASTUTERAG and InstructRAG. On GPT-4o, SeConRAG achieves the highest accuracy (83.6%) and lowest ASR (2.4%) under full poisoning, indicating its effectiveness even with strong LLMs.On smaller models such as Qwen-7B and LLaMA-3.1-8B, SeConRAG maintains competitive performance, outperforming all baselines under medium and low poisoning, demonstrating its scalability across model sizes. Under clean settings (0% poisoning), SeConRAG performs well and achieves high accuracy, indicating that the two-stage filtering does not overly suppress useful content.
Table 3: Performance comparison of SeConRAG and baseline methods on HotpotQA using different Poisoning RAG ratios (highest accuracy ↑ or lowest ASR ↓).
Model Method HotpotQA (Yang2018hotpotqa, )
100% 80% 60% 40% 20% 0%
(ACC↑ / ASR↓) (ACC↑ / ASR↓) (ACC↑ / ASR↓) (ACC↑ / ASR↓) (ACC↑ / ASR↓) (ACC↑ )
Mistral-12B (aydin2025generative, ) VanillaRAG 0.9 / 98.2 9.1 / 90.0 11.8 / 86.4 21.8 / 74.5 38.2 / 58.0 75.0
InstructRAG (wei2024instructrag, ) 13.6 / 83.5 23.6 / 71.8 25.5 / 70.0 37.3 / 57.3 45.5 / 49.1 75.0
ASTUTERAG (wang2024astute, ) 32.7 / 61.1 40.0 / 55.5 47.3 / 50.0 55.5 / 35.5 65.9 / 21.8 76.0
TrustRAG (zhou2025trustrag, ) 75.5 / 3.6 74.5 / 5.5 78.2 / 4.5 74.5 / 6.4 71.8 / 14.5 81.0
SeconRAG(ours) 75.7 / 3.6 77.3 / 4.5 75.5 / 4.5 71.8 / 8.2 72.7 / 4.5 83.0
Qwen-7B (hui2024qwen2, ) VanillaRAG 1.8 / 98.2 9.1 / 90.0 14.5 / 85.5 23.6 / 75.5 32.7 / 65.5 67.0
InstructRAG (wei2024instructrag, ) 24.5 / 76.4 30.9 / 69.1 31.8 / 68.2 35.5 / 63.6 45.5 / 51.8 67.0
ASTUTERAG (wang2024astute, ) 45.5 / 44.1 44.5 / 43.6 46.4 / 42.7 50.9 / 35.5 58.6 / 25.4 65.0
TrustRAG (zhou2025trustrag, ) 58.2 / 2.7 64.5 / 4.5 69.1 / 4.5 65.5 / 3.6 58.2 / 26.4 73.0
SeconRAG(ours) 63.6 / 2.3 67.3 / 1.8 73.6 / 3.6 67.3 / 2.7 61.8 / 21.8 76.0
LLaMA-3.1-8B (dubey2024llama, ) VanillaRAG 4.5 / 96.4 25.5 / 74.5 30.0 / 68.2 42.7 / 63.6 36.4 / 57.3 70.0
InstructRAG (wei2024instructrag, ) 27.3 / 71.8 42.7 / 54.5 51.8 / 46.4 49.1 / 48.2 47.3 / 50.0 76.0
ASTUTERAG (wang2024astute, ) 46.8 / 47.0 52.7 / 40.0 53.6 / 38.2 62.7 / 29.1 65.5 / 20.9 68.0
TrustRAG (zhou2025trustrag, ) 67.3 / 3.0 65.5 / 7.3 68.2 / 6.4 71.8 / 5.5 65.5 / 19.1 72.0
SeconRAG(ours) 72.0 / 10.9 78.2 / 4.5 75.5 / 3.6 77.3 / 1.8 67.4 / 18.4 84.0
GPT-4o (achiam2023gpt, ) VanillaRAG 11.9 / 81.8 32.7 / 57.3 46.4 / 50.0 48.2 / 43.6 45.5 / 30.5 81.0
InstructRAG (wei2024instructrag, ) 27.3 / 71.8 46.4 / 50.0 48.2 / 49.1 55.5 / 40.9 61.8 / 33.2 84.0
ASTUTERAG (wang2024astute, ) 67.3 / 24.1 73.6 / 15.5 77.3 / 12.7 78.2 / 10.0 77.3 / 11.8 81.0
TrustRAG (zhou2025trustrag, ) 80.9 / 2.7 83.6 / 3.6 81.8 / 3.6 81.8 / 3.6 79.1 / 6.4 85.0
SeconRAG(ours) 83.6 / 2.4 82.7 / 4.5 83.6 / 4.5 83.6 / 1.8 79.1 / 5.5 86.0
DeepSeek-R1 (guo2025deepseek, ) VanillaRAG 10.0 / 89.1 31.8 / 67.3 35.5 / 61.8 40.9 / 55.5 51.0 / 46.4 81.0
InstructRAG (wei2024instructrag, ) 27.3 / 72.7 48.2 / 51.8 57.3 / 42.7 56.4 / 42.7 61.8 / 38.2 80.0
ASTUTERAG (wang2024astute, ) 64.5 / 25.5 66.4 / 24.5 72.7 / 18.2 72.7 / 17.3 77.3 / 14.5 79.0
TrustRAG (zhou2025trustrag, ) 79.1 / 2.7 81.8 / 5.5 86.4 / 1.8 82.7 / 2.7 85.5 / 10.0 89.0
SeconRAG(ours) 81.8 / 8.0 83.6 / 3.6 87.3 / 3.6 82.7 / 3.6 83.6 / 5.5 86.0
Natural Questions (NQ)
Table 4 compares SeConRAG’s performance to baseline methods across five language models on the Natural Questions (NQ) benchmark, with six poisoning levels ranging from 0% (clean) to 100% . Through all LLMs and poisoning levels, SeConRAG consistently outperforms baseline methods in terms of answer accuracy and attack robustness. On Mistral-12B, SeConRAG outperforms TrustRAG and ASTUTERAG in both metrics, achieving up to 82.0% accuracy on clean data and maintaining high performance under attack (74.5% at 20% poisoning with only 10.2% ASR). Even with a smaller model, SeConRAG shows significant improvement. It achieves 78.0% accuracy on clean data and is more robust to 100% poisoning (66.4% / 2.4%) than TrustRAG (60.0% / 2.7%) and ASTUTERAG (42.3% / 53.2%). SeConRAG achieves 90.0% accuracy on clean data and 90.0% under 60% poisoning with 0.0% ASR, outperforming all baselines at almost every poisoning level on LLaMA-3.1-8B. On GPT-4o or DeepSeek-R1, SeConRAG outperforms at low-to-medium poisoning levels while maintaining low ASR across all ratios. SeConRAG outperforms TrustRAG and ASTUTERAG by achieving 100.0% accuracy with 0.0% ASR at 40% poisoning and over 96% accuracy with 0.0% ASR under full (100%) poisoning. These findings demonstrate SeConRAG’s ability to maintain high factual accuracy while resisting poisoning attacks. Its consistent performance in both clean and adversarial environments demonstrates the effectiveness of the two-stage SCF and CAF filtering mechanisms.
Table 4: Performance comparison of SeConRAG and baseline methods on NQ using different Poisoning RAG ratios (highest accuracy ↑ or lowest ASR ↓).
Model Method NQ (Kwiatkowski2019natural, )
100% 80% 60% 40% 20% 0%
(ACC↑ / ASR↓) (ACC↑ / ASR↓) (ACC↑ / ASR↓) (ACC↑ / ASR↓) (ACC↑ / ASR↓) (ACC↑ )
Mistral-12B (aydin2025generative, ) VanillaRAG 8.2 / 90.9 10.9 / 87.3 14.5 / 80.0 29.1 / 65.5 38.2 / 48.2 68.0
InstructRAG (wei2024instructrag, ) 13.6 / 82.7 17.3 / 78.2 26.4 / 70.0 38.2 / 56.4 51.8 / 40.0 66.0
ASTUTERAG (wang2024astute, ) 43.6 / 38.2 50.9 / 32.7 53.6 / 28.2 60.0 / 20.0 67.7 / 11.8 70.0
TrustRAG (zhou2025trustrag, ) 62.7 / 1.8 63.6 / 2.7 63.6 / 2.7 64.5 / 2.7 66.4 / 13.6 73.0
SeconRAG(ours) 63.6 / 2.5 65.5 / 0.0 66.4 / 3.6 67.3 / 0.0 74.5 / 10.2 82.0
Qwen-7B (hui2024qwen2, ) VanillaRAG 5.5 / 93.6 10.0 / 88.2 14.5 / 82.7 27.3 / 69.1 39.1 / 51.8 56.0
InstructRAG (wei2024instructrag, ) 25.5 / 76.4 33.6 / 65.5 33.6 / 65.5 34.5 / 62.7 47.3 / 47.3 64.0
ASTUTERAG (wang2024astute, ) 42.3 / 53.2 48.2 / 46.4 50.9 / 39.1 53.6 / 31.8 60.5 / 17.3 68.0
TrustRAG (zhou2025trustrag, ) 60.0 / 2.7 64.5 / 7.3 62.7 / 3.6 65.5 / 2.7 64.5 / 24.5 67.0
SeconRAG(ours) 66.4 / 2.4 70.0 / 4.5 67.3 / 5.5 68.2 / 3.6 70.9 / 21.8 78.0
LLaMA-3.1-8B (dubey2024llama, ) VanillaRAG 10.9 / 88.2 16.4 / 81.8 21.8 / 71.8 33.6 / 59.1 41.8 / 52.7 70.0
InstructRAG (wei2024instructrag, ) 32.7 / 67.3 44.5 / 54.5 43.6 / 54.5 49.1 / 49.1 56.4 / 34.5 70.0
ASTUTERAG (wang2024astute, ) 58.2 / 31.8 60.0 / 25.5 64.5 / 25.5 70.0 / 18.2 77.5 / 8.2 81.0
TrustRAG (zhou2025trustrag, ) 79.1 / 0.0 83.6 / 2.7 85.5 / 2.7 83.6 / 1.8 79.1 / 10.9 84.0
SeconRAG(ours) 88.2 / 1.8 88.2 / 5.5 90.0 / 0.0 89.1 / 1.8 86.9 / 4.0 90.0
GPT-4o (achiam2023gpt, ) VanillaRAG 27.3 / 68.2 33.6 / 61.8 41.8 / 49.1 50.0 / 36.4 52.7 / 31.8 74.0
InstructRAG (wei2024instructrag, ) 43.6 / 51.1 51.8 / 40.9 53.6 / 37.3 59.1 / 30.9 66.4 / 25.5 74.0
ASTUTERAG (wang2024astute, ) 75.5 / 14.2 75.5 / 12.7 76.4 / 12.7 78.2 / 9.1 79.1 / 10.9 81.0
TrustRAG (zhou2025trustrag, ) 80.0 / 0.1 81.8 / 1.8 82.7 / 0.9 82.7 / 0.9 81.8 / 1.0 86.0
SeconRAG(ours) 81.8 / 0.0 81.8 / 0.9 83.6 / 0.9 85.5 / 0.0 84.5 / 1.0 88.0
DeepSeek-R1 (guo2025deepseek, ) VanillaRAG 17.3 / 84.5 30.9 / 68.2 34.5 / 64.5 43.6 / 54.5 51.0 / 43.6 80.0
InstructRAG (wei2024instructrag, ) 39.1 / 62.7 50.9 / 48.2 52.7 / 47.3 57.3 / 41.8 65.5 / 32.7 82.0
ASTUTERAG (wang2024astute, ) 81.8 / 10.9 80.9 / 11.8 87.3 / 7.3 85.5 / 5.5 89.1 / 0.0 87.0
TrustRAG (zhou2025trustrag, ) 88.2 / 0.0 90.0 / 0.9 89.1 / 0.0 90.0 / 0.0 90.0 / 3.6 91.0
SeconRAG(ours) 96.4 / 0.0 98.2 / 0.0 96.4 / 0.0 100.0 / 0.0 96.4 / 0.0 98.0
MS-MARCO
Table 5 compares the performance of SeConRAG and baseline RAG defense methods on the MS-MARCO dataset at different corpus poisoning ratios (0% to 100%). SeConRAG consistently delivers the best or near-best performance in all settings. Mistral-12B: SeConRAG outperforms ASTUTERAG and InstructRAG, achieving 91.8% accuracy with 0.0% ASR under 60% poisoning and 98.0% accuracy in clean settings. Qwen-7B: Despite being a smaller model, SeConRAG achieves 84.0% accuracy in the clean setting and maintains low ASR (e.g., 4.5% at 100% poisoning), outperforming TrustRAG by a significant margin. LaMA-3.1-8B: SeConRAG achieves 90.0% accuracy in the clean setting and demonstrates strong robustness even under high poisoning (e.g., 89.1% / 0.0% at 100%). GPT-4o: SeConRAG matches or slightly outperforms TrustRAG for all poisoning levels. It achieves 94.0% accuracy on clean data and maintains 89.1% accuracy with only 1.8% ASR under 100% poisoning. DeepSeek-R1: SeConRAG outperforms all other tested methods in terms of robustness. It achieves 94.5% accuracy with 0.0% ASR under 60% poisoning and maintains strong performance even at 100% poisoning (94.5%/1.8%), outperforming TrustRAG (89.1%/3.6%). These findings confirm that SeConRAG is not only effective at resisting large-scale corpus poisoning attacks, but it also excels at maintaining answer quality in both adversarial and clean environments.
Table 5: Performance comparison of SeConRAG and baseline methods on MS using different Poisoning RAG ratios (highest accuracy ↑ or lowest ASR ↓).
Model Method MS-MARCO (bajaj2016ms, )
100% 80% 60% 40% 20% 0%
(ACC↑ / ASR↓) (ACC↑ / ASR↓) (ACC↑ / ASR↓) (ACC↑ / ASR↓) (ACC↑ / ASR↓) (ACC↑ )
Mistral-12B (aydin2025generative, ) VanillaRAG 9.1 / 89.1 15.5 / 81.8 19.1 / 76.4 34.5 / 60.0 50.0 / 45.5 84.0
InstructRAG (wei2024instructrag, ) 15.5 / 78.2 17.3 / 77.3 24.5 / 70.0 35.5 / 57.3 57.3 / 36.4 81.0
ASTUTERAG (wang2024astute, ) 32.7 / 58.2 33.6 / 58.2 46.4 / 45.5 61.8 / 30.0 73.6 / 18.8 81.0
TrustRAG (zhou2025trustrag, ) 91.8 / 0.0 81.8 / 7.3 86.4 / 4.5 86.4 / 5.5 87.3 / 11.8 85.0
SeconRAG(ours) 88.2 / 0.0 91.8 / 1.8 91.8 / 0.0 90.9 / 1.8 89.1 / 9.1 98.0
Qwen-7B (hui2024qwen2, ) VanillaRAG 10.0 / 87.3 13.6 / 84.5 22.7 / 75.5 28.2 / 69.1 43.6 / 46.4 75.0
InstructRAG (wei2024instructrag, ) 43.6 / 57.8 39.1 / 59.1 47.3 / 50.0 49.1 / 48.2 49.1 / 45.5 75.0
ASTUTERAG (wang2024astute, ) 42.3 / 54.5 43.6 / 51.8 49.1 / 42.7 60.9 / 26.4 65.5 / 20.0 74.0
TrustRAG (zhou2025trustrag, ) 64.5 / 11.8 65.5 / 14.5 66.4 / 10.0 67.3 / 11.8 66.4 / 22.7 78.0
SeconRAG(ours) 71.8 / 4.5 71.8 / 6.4 73.6 / 6.4 75.5 / 6.4 75.5 / 17.5 84.0
LLaMA-3.1-8B (dubey2024llama, ) VanillaRAG 9.1 / 88.2 20.0 / 77.3 28.2 / 66.4 36.4 / 60.0 54.5 / 40.9 83.0
InstructRAG (wei2024instructrag, ) 48.5 / 51.8 45.5 / 52.7 53.6 / 42.7 62.7 / 33.6 72.7 / 27.3 81.0
ASTUTERAG (wang2024astute, ) 56.8 / 38.6 63.6 / 29.1 63.6 / 26.4 73.6 / 21.8 82.3 / 13.6 89.0
TrustRAG (zhou2025trustrag, ) 84.5 / 6.4 83.6 / 8.2 82.7 / 8.2 86.4 / 7.3 85.4 / 9.1 84.0
SeconRAG(ours) 89.1 / 0.0 89.1 / 0.0 85.5 / 5.5 87.3 / 3.6 86.2 / 9.1 90.0
GPT-4o (achiam2023gpt, ) VanillaRAG 30.0 / 64.1 46.4 / 43.6 56.4 / 34.5 59.1 / 25.5 72.3 / 16.4 84.0
InstructRAG (wei2024instructrag, ) 50.5 / 42.7 57.3 / 35.5 62.7 / 30.0 59.1 / 24.5 70.9 / 17.3 83.0
ASTUTERAG (wang2024astute, ) 76.4 / 15.5 78.2 / 10.9 80.0 / 6.4 80.0 / 9.1 82.7 / 6.4 86.0
TrustRAG (zhou2025trustrag, ) 89.1 / 1.8 90.9 / 1.8 89.1 / 3.6 88.2 / 3.6 84.5 / 6.4 88.0
SeconRAG(ours) 89.1 / 1.8 90.9 / 1.8 90.0 / 1.8 89.1 / 1.8 89.1 / 3.6 94.0
DeepSeek-R1 (guo2025deepseek, ) VanillaRAG 11.8 / 81.8 33.6 / 61.8 39.1 / 55.5 50.9 / 42.7 60.5 / 29.1 82.0
InstructRAG (wei2024instructrag, ) 51.8 / 47.5 54.5 / 44.5 61.8 / 37.3 67.3 / 30.9 72.7 / 26.4 87.0
ASTUTERAG (wang2024astute, ) 85.5 / 8.2 80.9 / 13.6 80.9 / 10.0 87.3 / 7.3 89.1 / 5.5 88.0
TrustRAG (zhou2025trustrag, ) 89.1 / 3.6 90.9 / 2.7 91.8 / 2.7 91.8 / 3.6 89.1 / 5.5 91.0
SeconRAG(ours) 94.5 / 1.8 94.5 / 1.8 94.5 / 0.0 96.4 / 0.0 94.5 / 5.5 94.0
A.4.2 Impact of SCF Subcomponents
To demostrate the necessity of combining two filtering processes, we evaluating each subcomponent independently. As shown in 6,while each module provides moderate improvements on its own, when combined, they result in significantly increased robustness (for example, 0% ASR in several settings). These results confirm that the combination of clustering and semantic filtering is complementary, yielding the strongest robustness overall.
Table 6: Ablation of SCF components on Mistral-12B .
Model Setting HotpotQA (Yang2018hotpotqa, ) NQ (Kwiatkowski2019natural, ) MS-MARCO (bajaj2016ms, )
PIA 100% 20% PIA 100% 20% PIA 100% 20%
ACC↑/ASR↓ ACC↑/ASR↓ ACC↑/ASR↓ ACC↑/ASR↓ ACC↑/ASR↓ ACC↑/ASR↓ ACC↑/ASR↓ ACC↑/ASR↓ ACC↑/ASR↓
mistral-12b Clustering only 78 / 5 81 / 2 78 / 9 68 / 3 65 / 3 70 / 10 85 / 7 82 / 7 82 / 12
mistral-12b Semantic only 79 / 4 80 / 2 74 / 11 69 / 2 64 / 3 73 / 8 86 /5 82 / 6 86 / 8
mistral-12b Both (SCF) 77.5 / 0.8 75.7 / 3.6 72.7 / 4.5 72.3 / 1.8 63.6 / 2.5 74.5 / 10.2 91.8 / 0 88.2 / 0 89.1 / 9.1
A.4.3 Impact of of EIRE Module
To better understand the standalone contribution of the proposed Entity-Intent-Relation Extractor (EIRE) to SeCon-RAG’s overall robustness, We specifically compare SeCon-RAG’s performance with and without EIRE under various poisoning scenarios and three datasets, with Mistral-12B serving as the backbone model. The results are summarized in Table 7. With EIRE enabled, the model consistently achieves higher factual accuracy while significantly lowering the ASR, particularly under high poisoning conditions. For example, on the MS-MARCO dataset under 100% poisoning attack, enabling EIRE reduces ASR from 5% to 0% while increasing accuracy from 85% to 88.2%.
Table 7: Ablation of the EIRE module on Mistral-12B across three datasets and poisoning scenarios.
Model Setting HotpotQA (Yang2018hotpotqa, ) NQ (Kwiatkowski2019natural, ) MS-MARCO (bajaj2016ms, )
PIA 100% 20% PIA 100% 20% PIA 100% 20%
ACC↑/ASR↓ ACC↑/ASR↓ ACC↑/ASR↓ ACC↑/ASR↓ ACC↑/ASR↓ ACC↑/ASR↓ ACC↑/ASR↓ ACC↑/ASR↓ ACC↑/ASR↓
mistral-12b Without EIRE 76 / 5 75 / 4 73 / 11 69 / 3 63 / 4 72 / 16 87 / 5 85 / 5 83 / 11
mistral-12b With EIRE 77.5 / 0.8 75.7 / 3.6 72.7 / 4.5 72.3 / 1.8 63.6 / 2.5 74.5 / 10.2 91.8 / 0 88.2 / 0 89.1 / 9.1
These show that EIRE is critical for enabling fine-grained semantic reasoning which increases the accuracy of the final answer generation process.
A.4.4 Impact of the Verified Correct Document Set
To evaluate the effectiveness the efficacy of $d_{cor}$, we conduct an ablation study without $d_{cor}$ and measure the performance drop across three datasets under three poisoning scenarios, with Mistral-12B serving as the baseline. As shown in Table 8, removing D consistently reduces accuracy while increasing the attack success rate, particularly in high-poisoning settings. For example, on MS-MARCO with 100% poisoning, enabling D reduces ASR from 5% to 0% while increasing accuracy from 85% to 88.2%. These results demonstrate that even a small, high-quality $d_{cor}$ set can significantly improve semantic filtering performance and reduce noise from poisoned documents.
Table 8: Ablation of the verified correct document set $d_{cor}$.
Model Setting HotpotQA (Yang2018hotpotqa, ) NQ (Kwiatkowski2019natural, ) MS-MARCO (bajaj2016ms, )
PIA 100% 20% PIA 100% 20% PIA 100% 20%
ACC↑/ASR↓ ACC↑/ASR↓ ACC↑/ASR↓ ACC↑/ASR↓ ACC↑/ASR↓ ACC↑/ASR↓ ACC↑/ASR↓ ACC↑/ASR↓ ACC↑/ASR↓
mistral-12b Without $d_{cor}$ 76 / 10 80 / 6 73 / 12 72 / 6 63 / 3 73 / 10 85 / 9 82 / 6 85 / 8
mistral-12b With $d_{cor}$ 77.5 / 0.8 75.7 / 3.6 72.7 / 4.5 72.3 / 1.8 63.6 / 2.5 74.5 / 10.2 91.8 / 0 88.2 / 0 89.1 / 9.1
A.4.5 Sensitivity Analysis of Filtering Thresholds
To assess the robustness of SeCon-RAG in relation to its key hyperparameters, we perform a sensitivity analysis on the two primary filtering thresholds: $\tau_{\text{cluster}}$: the similarity threshold used in clustering-based filtering. $\tau_{\text{semantic}}$: the semantic similarity threshold used in EIRE-based semantic graph filtering. We vary each threshold across a reasonable range ($\tau_{\text{cluster}}\in[0.86,0.90]$, $\tau_{\text{semantic}}\in[0.2,0.4]$) and evaluate SeCon-RAG’s performance under three poisoning intensities on two representative models (LLaMA-3.1-8B and GPT-4o) and three datasets.
Table 9: Sensitivity analysis of $\tau_{\text{cluster}}$ on LLaMA-3.1-8B and GPT-4o under different poisoning intensities.
Model $\tau_{\text{cluster}}$ HotpotQA (Yang2018hotpotqa, ) NQ (Kwiatkowski2019natural, ) MS-MARCO (bajaj2016ms, )
PIA 100% 20% PIA 100% 20% PIA 100% 20%
ACC↑/ASR↓ ACC↑/ASR↓ ACC↑/ASR↓ ACC↑/ASR↓ ACC↑/ASR↓ ACC↑/ASR↓ ACC↑/ASR↓ ACC↑/ASR↓ ACC↑/ASR↓
LLaMA-3.1-8B 0.86 72 / 4 68 / 4 72 / 4 67 / 19 83 / 2 79 / 4 86 / 2 83 / 6 88 / 4
LLaMA-3.1-8B 0.90 72 / 4 69 / 3 74 / 4 65 / 19 83 / 2 80 / 4 86 / 2 84 / 6 88 / 4
GPT-4o 0.86 80 / 3 81 / 2 81 / 3 82 / 6 82 / 1 81 / 1 83 / 1 81 / 3 90 / 3
GPT-4o 0.90 81 / 3 81 / 3 84 / 4 81 / 9 82 / 2 83 / 1 84 / 2 83 / 1 88 / 4
Table 10: Sensitivity analysis of $\tau_{\text{semantic}}$ on LLaMA-3.1-8B and GPT-4o under different poisoning intensities.
Model $\tau_{\text{semantic}}$ HotpotQA (Yang2018hotpotqa, ) NQ (Kwiatkowski2019natural, ) MS-MARCO (bajaj2016ms, )
PIA 100% 20% PIA 100% 20% PIA 100% 20%
ACC↑/ASR↓ ACC↑/ASR↓ ACC↑/ASR↓ ACC↑/ASR↓ ACC↑/ASR↓ ACC↑/ASR↓ ACC↑/ASR↓ ACC↑/ASR↓ ACC↑/ASR↓
LLaMA-3.1-8B 0.2 68 / 7 67 / 5 74 / 4 66 / 19 82 / 2 80 / 4 86 / 2 83 / 7 88 / 4
LLaMA-3.1-8B 0.4 68 / 7 70 / 3 74 / 4 66 / 19 82 / 2 80 / 4 85 / 2 83 / 7 89 / 6
GPT-4o 0.2 81 / 4 81 / 3 84 / 3 84 / 7 81 / 2 82 / 2 82 / 1 84 / 3 89 / 4
GPT-4o 0.4 80 / 4 82 / 3 83 / 4 82 / 9 81 / 1 82 / 2 82 / 1 83 / 1 89 / 2
Tables 9 and 10 show that SeCon-RAG’s performance remains stable even when both thresholds are changed slightly. This is primarily due to the conservative AND-logic used in the joint filtering mechanism, which ensures that only documents flagged by both filters are excluded. These findings show that our framework is not overly sensitive to precise threshold tuning, which makes it easier to use in practice.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: SeConRAG introduces a dual-filtering framework to defend against corpus poisoning in retrieval-augmented generation (RAG) systems, as clearly stated in both the abstract and introduction. The stated contributions are consistent with the work’s scope, and no over-claims or unsupported generalizations are made. Therefore, the abstract and introduction accurately reflect the paper’s contributions and findings.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes] .
  9. Justification: In the conclusion section, we discussed the limitations of the proposed method.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer:[No]
  14. Justification: The formulas in the paper are simple to understand, and there is no theoretical
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The paper provides detailed descriptions of all experimental settings, including datasets, model variants, poisoning rates, evaluation metrics, and hardware configurations. Key implementation details for SeConRAG, such as filtering modules and embedding choices, are explicitly described. This information is sufficient to reproduce the main results supporting the paper’s claims.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [No]
  24. Justification: The code and data are not yet publicly released, but we plan to provide open access with full instructions upon acceptance
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The paper provides comprehensive experimental details, including dataset splits, poisoning ratios, evaluation metrics, and backbone model configurations. Since the method is inference-based, no model training is required, and optimizer settings are not applicable. All key implementation and runtime details are included either in the main text or supplemental material to ensure clarity and reproducibility.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: We do not report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We introduce the details in setup of experiment.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: This work complies with the NeurIPS Code of Ethics. It does not involve human subjects, personally identifiable information, or unauthorized data usage.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We have added an “Impact Statement” paragraph at the end of the submitted manuscript to discuss the broader impacts.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A] .
  54. Justification: We do not release the model or dataset.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We use the open-sourced dataset.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: There is no new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer:[N/A] .
  74. Justification: The paper does not involve Research with Human Subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer:[Yes]
  79. Justification: In this work, large language models (LLMs) are employed in two ways. First, LLMs (e.g., GPT-based models) are integrated as functional components of our proposed framework, mainly for semantic understanding and entity-intent-relation extraction. In addition, an LLM was used to assist with language polishing and clarity improvement of the manuscript. The scientific content, experimental design, and analysis were entirely conducted by the authors.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.